OpenStudy (lovelyharmonics):

Write the expression as either the sine, cosine, or tangent of a single angle. cosine of pi divided by five times cosine of pi divided by seven plus sine of pi divided by five times sine of pi divided by seven.

OpenStudy (campbell_st):

perhaps you could use the draw option to write it

OpenStudy (lovelyharmonics):

\[\cos \left( \frac{ π }{ 5 } \right)\cos \left( \frac{ π }{ 7 } \right) + \sin \left( \frac{ π }{ 5 } \right)\sin \left( \frac{ π }{ 7 } \right)\]

OpenStudy (lovelyharmonics):

would you start by multiplying the first two together and the second two together?

OpenStudy (anonymous):

probably, been a while since i have worked with angles like this

OpenStudy (campbell_st):

ok... so this looks like the cos identity for the difference of 2 angles \[\cos(A - B) = \cos(A)\cos(B) + \sin(A)\sin(B)\] hopefully this helps

OpenStudy (campbell_st):

so you know the angles... just substitute both angles into the left side of the equation and simplify

OpenStudy (lovelyharmonics):

\[\cos \left( \frac{ \pi }{ 5 } - \frac{ \pi }{ 7 }\right)\] so its this? @campbell_st

OpenStudy (lovelyharmonics):

that p is suppose to be pi....

OpenStudy (campbell_st):

thats it... just simplfy the 2 fractions... its like 1/5 - 1/7 but you need pi in the numerator

OpenStudy (campbell_st):

or \[\pi \times (\frac{1}{5} - \frac{1}{7}) = \]

OpenStudy (lovelyharmonics):

\[\cos \left( \frac{ 7\pi }{ 35 } - \frac{ 5\pi }{ 35 } \right)\]

OpenStudy (campbell_st):

yep... just simplify that for the answer

OpenStudy (lovelyharmonics):

\[\cos \left( \frac{ 2\pi }{ 35 } \right)\] this cant be simplified any further can it?

OpenStudy (anonymous):

no thats as far as you can simplify unless you solve

OpenStudy (lovelyharmonics):

okay thank you

OpenStudy (anonymous):

yw